# 1]Independent Researcher

### v1

@@ Abstract
This paper investigates data-driven solutions and parameter discovery to (2+ 1)-dimensional coupled nonlinear Schrödinger equations with variable coefficients (VC-CNLSEs), which describe transverse effects in optical fiber systems under perturbed dispersion and nonlinearity. By setting different forms of perturbation coefficients, we aim to recover the dark and anti-dark one- and two-soliton structures by employing an enhanced physics-based deep neural network algorithm, namely a physics-informed neural network (PINN). The enhanced PINN algorithm leverages the locally adaptive activation function mechanism to improve convergence speed and accuracy. In the lack of data acquisition, the PINN algorithms will enhance the capability of the neural networks by incorporating physical information into the training phase. We demonstrate that applying PINN algorithms to (2+ 1)-dimensional VC-CNLSEs requires distinct distributions of physical information. To address this, we propose a region-specific weighted loss function with the help of residual-based adaptive refinement strategy. In the meantime, we perform data-driven parameter discovery for the model equation, classified into two categories: constant coefficient discovery and variable coefficient discovery. For the former, we aim to predict the cross-phase modulation constant coefficient under varying noise intensities using enhanced PINN with a single neural network. For the latter, we employ a dual-network strategy to predict the dynamic behavior of the dispersion and non-linearity perturbation functions. Our study demonstrates that the proposed framework holds significant potential for studying high-dimensional and complex solitonic dynamics in optical fiber systems.
@@
Physics-informed neural networks, Dual-network, (2+ 1)-dimensional coupled non-linear Schrödinger equations, Data-driven solitons, Parameter discovery
@@ 1 Introduction
In the field of optical fiber communication, nonlinear differential equations are commonly used to model the behavior of light propagation through the fiber. These equations take into account factors such as dispersion, nonlinearity, and attenuation to accurately predict the signal transmission in optical fibers. Nonlinear differential equations are essential for understanding the complex interactions between light and the fiber medium, allowing researchers to optimize signal transmission and minimize data loss. By incorporating these factors into mathematical models, engineers can design more efficient and reliable optical communication systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. One of the most important equations used in this context is the nonlinear Schrödinger equation (NLSE). The NLSE is a fundamental partial differential equation that describes the evolution of the slowly varying envelope of a pulse of light in an optical fiber, considering the effects of both dispersion and nonlinearity. Specifically, it captures the balance between the dispersive effects, which tend to spread the pulse in time, and the nonlinear effects, such as self-phase modulation, which can lead to pulse shaping or even soliton formation, where the pulse maintains its shape over long distances. The NLSE and its extensions have been widely studied in many scientific disciplines, including nonlinear optics [15, 16, 17], fluid dynamics [18], quantum mechanics [19, 20], and plasma physics [21, 22]. By solving the NLSE, researchers can predict and manipulate the behavior of light pulses in optical fibers, enabling advancements in telecommunications and optical signal processing. The standard NLSE can be formulated as
$i\psi_{t}+\alpha\psi_{xx}+\beta\lvert\psi\rvert^{2}\psi=0,$
where $\psi=\psi(x,t)$ represents a complex amplitude of two variables $x$ and $t$, $\alpha$ and $\beta$ are real constants, which affect the waveform broadening and the oscillation during light wave transmission, respectively. $i$ denotes the imaginary units and the subscript $x$ $(t)$ represents the partial derivative with respect to $x$ $(t)$. The second term $\psi_{xx}$ represents the group velocity dispersion, while the third term denotes the nonlinearity. In the past few years, many numerical methods have been developed to study the dynamic behavior of NLSE, such as inverse scattering transform, Hirota method, Sardar sub-equation technique, and so on [23, 24, 25, 26, 27, 28, 29, 30]. Many extensions to standard NLSE have been introduced to model various physical phenomena for a more accurate representation of complex systems. As one of the most effective integrable models, derivative NLSE is studied in the field of plasma physics to model Alfvén waves propagation [31, 32]. The perturbed NLSE describes the pulse propagation of high-order dispersion in an optical fiber, when realistic fibers are considered [33]. An integrable nonlocal NLSE has been introduced in [34], which is parity-time(PT)-symmetric. When considering the propagation of waveforms through multiple fields at different polarizations and frequencies, the coupled NLSE (CNLSE) can be formulated as [35, 36, 37]
$i\psi_{1,t}$ $+\psi_{1,xx}+\beta(\lvert\psi_{1}\rvert^{2}+\tau\lvert\psi_{2}\rvert^{2})\psi_{1}=0,$
$i\psi_{2,t}$ $+\psi_{2,xx}+\beta(\tau\lvert\psi_{1}\rvert^{2}+\lvert\psi_{2}\rvert^{2})\psi_{2}=0,$ (1)
where $\psi_{1}=\psi_{1}(x,t)$ and $\psi_{2}=\psi_{2}(x,t)$ are the amplitudes of wave propagation through a two-mode optical fiber, the value of the parameter $\beta=\pm 1$ is used to differentiate between the self-focusing and self-defocusing Kerr nonlinearity and the range of the parameter $\tau$ varies, $\tau>2/3$ and $\tau<7$, regarding the Kerr-shape electron nonlinearity and the nonlinearity resulting from molecular orientation changes [38, 39]. In [40], general $N$-dark-dark solitons for system (1) are obtained by applying the KP-hierarchy reduction method. Localized wave solutions including dark-dark and bright-dark solitons, breather-breather solutions, and different types of new vector rogue wave solutions as well as their interactions are studied for defocusing version of system (1) through the application of the Darboux transformations [41]. By virtue of the Hirota method, a varity of optical solitons are obtained for high-dimensional generalized CNLSE in the field of optical fiber communication [42, 43, 44].
In the field of optical fiber communication, studying the dynamic behavior of optical solitons is crucial for understanding pulse propagation, enhancing data transmission rates, and mitigating signal degradation over long distances. In the past decades, many formations of solitons have been observed in different optical fiber systems, including bright solitons, dark solitons, vector solitons, and so on. The reason for obtaining different formations of optical solitons for NLSE lies in the varying balance between dispersion and nonlinearity within the optical medium, as well as the initial conditions and parameters such as pulse shape, power, and wavelength. These different types of solitons exhibit unique properties that can be harnessed for various applications in optical communication systems [45, 46, 47]. The optical soliton solutions can maintain the shape and speed of optical pulses over long distances, effectively countering the effects of dispersion and nonlinearity in the fiber, and thus enabling high-performance and reliable data transmission.
While lower-dimensional versions of CNLSE describe the evolution of wave packets in a single spatial direction, higher-dimensional CNLS equations can model more complex physical phenomena in higher-dimensional scenarios with a broader range of behaviors and applications. Due to the presence of additional spatial variables, it is possible to model wave propagation in larger spaces rather than along a single line. The higher-dimensional CNLSE is particularly relevant for understanding phenomena like the formation and interaction of vortex solitons, two-dimensional localized structures, and pattern formation in nonlinear media. In contrast, the (1+ 1)-dimensional CNLSE is limited to simpler solitary wave solutions, such as bright and dark solitons, that are confined to one spatial direction. Due to the complexity of modeling with higher-dimensional versions, advanced numerical techniques and approximations are required for analysis. The (2+ 1)-dimensional CNLSE can be constructed as
$i\psi_{1,t}$ $+\psi_{1,xx}+\psi_{1,yy}+\beta(\lvert\psi_{1}\rvert^{2}+\tau\lvert\psi_{2}\rvert^{2})\psi_{1}=0,$
$i\psi_{2,t}$ $+\psi_{2,xx}+\psi_{2,yy}+\beta(\tau\lvert\psi_{1}\rvert^{2}+\lvert\psi_{2}\rvert^{2})\psi_{2}=0,$ (2)
where $\psi_{1}=\psi_{1}(x,y,t)$ and $\psi_{2}=\psi_{2}(x,y,t)$ describe the amplitudes of two circularly polarized waves. The solitonic structures of system (1) and its extensions have been discovered using a variety of numerical techniques [48, 49, 50, 51, 52]. In optical fiber communication, many mathematical models have been studied the evolution of complex wave envelopes in homogeneous media, where the related parameters remain fixed throughout the medium. However, many real-world systems exhibit spatial or temporal variations in their nature, such as inhomogeneous plasmas and varying refractive indices. To address these variations, the variable coefficient NLSE is introduced, where the coefficients related to dispersion and nonlinearity are allowed to vary with respect to spatial and/or temporal variables [53, 54, 55, 56]. Considering the local defects and damages in the optical fibers, the variable coefficient (2+ 1)-dimensional CNLSE (VC-CNLSE) with perturbed dispersion and nonlinearity can be constructed as
@@
$i\psi_{2,t}$ $+\alpha(t)(\psi_{2,xx}+\psi_{2,yy})+\beta(t)(\tau\lvert\psi_{1}\rvert^{2}+\lvert\psi_{2}\rvert^{2})\psi_{2}=0.$ (3)
@@
By choosing different forms of perturbation functions $\alpha(t)$ and $\beta(t)$, one can obtain various optical solitons by controlling factors such as intensities, velocities, and accelerations. By virtue of the Hirota method, linear, parabolic, and quasi-periodic dark solitons are discovered in [57]. Abundant dark and anti-dark soliton structures, as well as their collisions are studied in [58] using the Hirota method. By utilizing the developed Hirota bilinear method, one-soliton, two-soliton, three-soliton, and breather like bright-dark solitons are observed in [59].
In the era that data analysis has become increasingly important, machine learning algorithms have emerged as powerful tools for extracting insights and making predictions from large datasets. These algorithms can automatically learn and improve from experience without being explicitly programmed, allowing for more efficient and accurate analysis of complex data patterns. Machine learning algorithms have been applied in many scientific disciplines, including natural language processing [60], image recognition [61], bioinformatics [62, 63], and climate modeling [64]. These diverse applications underscore the versatility of machine learning as a tool for solving complex data-driven problems across various domains. Deep learning is a machine learning technique that uses the universal approximation capability of neural networks to extract and learn hidden patterns within available data. These neural networks, composed of multiple layers of interconnected nodes, allow deep learning models to automatically identify and represent complex features in data. The success of deep learning can be attributed to its ability to process vast amounts of data and uncover intricate relationships that may not be apparent through traditional methods [65, 66]. By leveraging large datasets, advanced architectures, and powerful computational resources, deep learning has revolutionized numerous fields. The performance and accuracy of deep learning models are directly related to the amount of available data to be analyzed. In the lack of sufficient data acquisition, machine learning algorithms face challenges that can lead to poor performance and robustness, with no guarantee of convergence. In these scenarios, using the prior knowledge of the problem at hand in the modeling process can lead the solution space into the admissible one. This prior information can be represented through various forms, such as domain-specific constraints, expert knowledge, or physics-based models. This approach is particularly useful in fields where data acquisition is expensive to obtain, ensuring that the model remains reliable and effective. Recently, we have witnessed the successful application of deep learning methods in the field of solving differential equations, leading to the formation of a new branch of studies, namely scientific machine learning.
As a physics-based methodology, physics-informed neural networks (PINNs) have been effectively applied in various scientific scenarios such as fluid dynamics, material science, and medical imaging. PINNs combine the power of neural networks with the physical principles governing a system to accurately predict outcomes and simulate complex phenomena [67]. In contrast to traditional numerical methods, PINNs enhance the approximation capability of deep neural networks by incorporating physical laws, described by differential equations, into the training phase. These physical laws can act as a regularization to constraint the solution space of neural networks. Training these neural networks involves solving a minimization problem, where a multi-objective loss function – combining physics-informed and data-driven losses – is optimized. This approach results in accurate predictions while ensuring consistency with the underlying physical principles. By considering the spatial and/or temporal coordinates as input data, PINNs leverage the Automatic Differentiation (AD) technique [68], enabling the efficient computation of derivatives with respect to the input variables, which are necessary to enforce the governing physical laws. Then, a multi-objective loss function can be constructed using the residual forms of the boundary and/or initial conditions, as well as the residuals related to the differential equations. This loss function is optimized during training, allowing the neural network to approximate solutions that not only fit the data but also adhere to the underlying physical principles. As a result, PINNs are able to provide accurate and physically consistent solutions even in cases where traditional methods may struggle, particularly when data is scarce or noisy [69]. While PINNs offer innovative approaches and flexibility, there are many challenges and limitations that need to be addressed [70, 71]. Some of the advantages of PINNs methodologies are that they are capable of handling problems with irregular geometries and boundary conditions, highly flexible in dealing with nonlinear problems, and able to solve inverse problems where traditional methods may struggle. Additionally, they are easy to code, with no significant differences between the coding for forward and inverse problem solving.
@@
In recent years, numerous studies have been reported on the applications of PINNs in the field of optical fiber communication. In [72], an adaptive residual points PINN scheme is introduced for vector soliton simulation and parameter discovery of general CNLSE. Data-driven vector soliton solutions and parameter discovery for the coupled mixed derivative NLSE are studied in [73], where a PINN method with a twin subnet strategy is used. In [74], an improved PINN algorithm is introduced to simulate both vector degenerate and nondegenerate solitons of the coupled nonlocal NLSE and to estimate the associated equation parameters under varying noise intensities. A PINN scheme with a locally adaptive activation function is employed in [75] for data-driven localized wave solutions of the derivative NLSE. In [76], the standard PINNs is used for prediction of soliton dynamics and estimation of model parameters for (2+ 1)-dimensional NLSE. In [77], the authors employed an improved PINN algorithm for solving forward and inverse problems, where the model equation was the Hirota equation with variable coefficients. They applied the proposed algorithm for data-driven discovery of the constant coefficients and the linear function variable coefficients in the model equation.
@@
Most of the studies mentioned above have focused on simulating the propagation of waveforms, modeled by the NLSE and its extensions, in a homogeneous environment. Traditional numerical methods face challenges in simulating high-dimensional problems due to the curse of dimensionality, which causes an exponential increase in computational cost and complexity as the number of dimensions grows. When dealing with strongly nonlinear problems, these methods may encounter difficulties such as convergence issues and numerical instability. This is particularly problematic when the coupled equations involve interactions across multiple scales or when the solution exhibits sharp gradients or localized structures like solitons. These limitations have driven the development of alternative approaches, such as PINNs, to address the shortcomings of traditional methods in modern applications. In this paper, an enhanced PINN methology is introduced for the prediction of soliton structures to (2+ 1)-dimensional CNLSE (1). To the best of our knowledge, the application of PINNs to this class of equations has not been reported before. The proposed approach leverages a neuron-wise locally adaptive activation function to accelerate the convergence rate and improve the accuracy of standard PINN [78]. By combining the proposed approach with the residual-based adaptive refinement strategy (RAR) [79], we demonstrate that each residual loss constructed from the model Eq. (1) requires a different distribution of residual points within the computational domain. The performance and efficiency of the proposed framework will be evaluated in predicting various soliton dynamics through the lens of $L^{2}$ relative error. By considering different forms of dispersion and nonlinearity coefficients, vector parabolic dark one-soliton, vector m-shaped anti-dark one-soliton, and dark two-soliton are predicted. We also apply the proposed framework to the parameter discovery of the model Eq. (1). First, we apply the introduced PINN, equipped with a single neural network, for constant parameter discovery under different noise intensities. This PINN model not only predicts the solution of the model equation but also estimates the constant parameter as another trainable network parameter during training. Second, we employ a dual-network strategy for the proposed PINN scheme, where one neural network predicts the solution dynamics, while the other captures the behavior of the dispersion and nonlinearity variable coefficients. The obtained results indicate that the proposed framework is promising for exploring waveform propagation in the optical fiber communications landscape.
@@
This is how the remainder of the paper is organized. In Sect. 2, we introduce the general framework of PINN algorithms for the general (2+ 1)-dimensional variable coefficient coupled nonlinear equations and discuss the related mechanisms. We also introduce our enhanced PINN framework in this section. We investigate the application of the proposed enhanced PINN algorithm for data-driven vector dark and anti-dark one- and two-soliton solutions to the (2+ 1)-dimensional VC-CNLSE. We illustratively show that our proposed algorithm accurately captures the solitonic dynamics. We also perform data-driven parameter discovery for the system of model equations to estimate the value of constant coefficient as well as the variable dispersion and nonlinearity coefficients. A conlusion is made in Sec. 4, where we summarize the effectiveness of the enhanced PINN framework and discuss its potential.

### v2

# Enhanced PINNs for data-driven solitons and parameter discovery for (2+ 1)-dimensional coupled nonlinear Schrödinger systems

@@ Abstract
This paper investigates data-driven solutions and parameter discovery to (2+ 1)-dimensional coupled nonlinear Schrödinger equations with variable coefficients (VC-CNLSEs), which describe transverse effects in optical fiber systems under perturbed dispersion and nonlinearity. By setting different forms of perturbation coefficients, we aim to recover the dark and anti-dark one- and two-soliton structures by employing an enhanced physics-based deep neural network algorithm, namely a physics-informed neural network (PINN). The enhanced PINN algorithm leverages the locally adaptive activation function mechanism to improve convergence speed and accuracy. In the lack of data acquisition, the PINN algorithms will enhance the capability of the neural networks by incorporating physical information into the training phase. We demonstrate that applying PINN algorithms to (2+ 1)-dimensional VC-CNLSEs requires distinct distributions of physical information. To address this, we propose a region-specific weighted loss function with the help of residual-based adaptive refinement strategy. In the meantime, we perform data-driven parameter discovery for the model equation, classified into two categories: constant coefficient discovery and variable coefficient discovery. For the former, we aim to predict the cross-phase modulation constant coefficient under varying noise intensities using enhanced PINN with a single neural network. For the latter, we employ a dual-network strategy to predict the dynamic behavior of the dispersion and nonlinearity perturbation functions. Our study demonstrates that the proposed framework holds significant potential for studying high-dimensional and complex solitonic dynamics in optical fiber systems.
@@
Physics-informed neural networks, Dual-network, (2+ 1)-dimensional coupled nonlinear Schrödinger equations, Data-driven solitons, Parameter discovery.
[…]
@@ 1 Introduction
In optical fiber communication, nonlinear differential equations are widely used to model light propagation behavior, accounting for dispersion, nonlinearity, and attenuation to accurately predict signal transmission. These equations help researchers understand complex light-fiber interactions, enabling the design of efficient and reliable communication systems. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. In this context, the nonlinear Schrödinger equation (NLSE) and its extensions play a crucial role in understanding pulse dynamics and soliton formation in optical fibers. These equations have been widely studied in many scientific disciplines, including nonlinear optics [15, 16, 17], fluid dynamics [18], quantum mechanics [19, 20], and plasma physics [21, 22].
To model more realistic systems, various extensions of the standard NLSE have been introduced, including the coupled NLSE (CNLSE), which describes pulse propagation in multi-mode or birefringent fibers. Furthermore, higher-dimensional versions, such as the (2+ 1)-dimensional CNLSE, have been employed to study complex spatio-temporal structures like vortex solitons and two-dimensional localized waves. However, real-world optical fibers often exhibit spatial or temporal variations due to inhomogeneities, imperfections, or external perturbations, which cannot be captured by constant-coefficient models. To address this, the variable-coefficient (2+ 1)-dimensional coupled NLSE (VC-CNLSE) has been proposed, in which the dispersion and nonlinearity coefficients vary with respect to time or space [24, 25, 26, 27]. The (2+ 1)-dimensional VC-CNLSE can be formulated as:
@@
$i\psi_{2,t}$ $+\alpha(t)(\psi_{2,xx}+\psi_{2,yy})+\beta(t)(\tau\lvert\psi_{1}\rvert^{2}+\lvert\psi_{2}\rvert^{2})\psi_{2}=0,$ (1)
@@
where $i$ denotes the imaginary unit, $\psi_{1}(x,y,t)$ and $\psi_{2}(x,y,t)$ are both complex functions, representing the amplitudes of wave propagation through a two-mode optical fiber, and the subscripts $x$, $y$, and $t$ denote the partial derivatives with respect to these variables. The parameter $\tau$ varies, $\tau>2/3$ and $\tau<7$, regarding the Kerr-shape electron nonlinearity and the nonlinearity resulting from molecular orientation changes. By choosing different forms of perturbation functions $\alpha(t)$ and $\beta(t)$, one can obtain various optical soliton solutions by controlling factors such as intensity, velocity, and acceleration. Many studies have investigated solutions to the VC-CNLSE (1) in recent years. By virtue of the Hirota method, linear, parabolic, and quasi-periodic dark solitons were reported in [28]. Abundant dark and anti-dark soliton structures, as well as their collisions, are studied in [29] using the Hirota method. By utilizing the developed Hirota bilinear method, one-soliton, two-soliton, three-soliton, and breather-like bright-dark solitons are observed in [30].
With the growing importance of data analysis, machine learning has emerged as a powerful tool for extracting insights and making predictions from large datasets, with applications in natural language processing [31], image recognition [32], bioinformatics [33, 34], and climate modeling [35]. Deep learning, a machine learning paradigm that leverages the universal approximation capability of neural networks, automatically learns complex features from data through its multi-layered architecture. Its success lies in processing vast amounts of data to uncover intricate patterns that are not apparent to traditional methods [36, 37]. However, model performance depends heavily on data availability; insufficient data can lead to poor robustness and lack of convergence. Incorporating prior knowledge—such as domain constraints, expert input, or physics-based models—can guide solutions toward admissible spaces, especially where data acquisition is costly. Recently, deep learning has been successfully applied to solving differential equations, giving rise to the emerging field of scientific machine learning.
Physics-informed neural networks (PINNs) combine neural networks with physical laws to predict and simulate complex phenomena accurately [38]. Unlike traditional numerical methods, PINNs embed governing differential equations into the training process as a form of regularization, thereby enhancing their approximation capabilities. They treat spatial and temporal coordinates as inputs and employ automatic differentiation (AD) [39] for efficient derivative computation, enforcing physical constraints through a multi-objective loss that integrates both data-driven and physics-informed residuals. This framework ensures that predictions remain consistent with underlying physical principles. PINNs have found applications in fluid dynamics, materials science, and medical imaging, offering flexibility for nonlinear, irregular, or inverse problems where traditional methods often struggle. Nevertheless, challenges and limitations remain [40, 41]. Moreover, PINNs are relatively straightforward to implement, with similar coding structures for both forward and inverse problems.
@@
In recent years, numerous studies have been reported on the applications of PINNs in the field of optical fiber communication. In [42], an adaptive residual points PINN scheme is introduced for vector soliton simulation and parameter discovery of general CNLSE. Data-driven vector soliton solutions and parameter discovery for the coupled mixed derivative NLSE are studied in [43], where a PINN method with a twin subnet strategy is used. In [44], an improved PINN algorithm is introduced to simulate both vector degenerate and nondegenerate solitons of the coupled nonlocal NLSE and to estimate the associated equation parameters under varying noise intensities. A PINN scheme with a locally adaptive activation function is employed in [45] for data-driven localized wave solutions of the derivative NLSE. In [46], the standard PINNs is used for prediction of soliton dynamics and estimation of model parameters for (2+ 1)-dimensional NLSE. In [47], the authors employed an improved PINN algorithm for solving forward and inverse problems, where the model equation was the Hirota equation with variable coefficients. They applied the proposed algorithm for data-driven discovery of the constant coefficients and the linear function variable coefficients in the model equation. In [48], the authors extend PINNs to model stationary and non-stationary solitons of $1D$ and $2D$ saturable nonlinear Schrödinger equations (SNLSEs) with PT-symmetric potentials. They also propose a modified PINNs (mPINNs) to directly identify potential functions from solution data, investigate inverse problems, and compare network structures. A residual-unit-based gradient-enhanced PINN (R-gPINN) for solving forward and inverse problems of variable-coefficient PDEs is proposed in [49]. The authors introduced residual units to address gradient vanishing and network degradation, and incorporates gradient terms of variable coefficients to better enforce physical constraints. Through numerical experiments on several equations (Burgers, KdV, Sine-Gordon, and KP), they show that R-gPINN significantly improves the accuracy of both solution prediction and coefficient discovery.
@@
Most studies above focus on simulating waveform propagation, modeled by the NLSE and its extensions, in homogeneous environments. Traditional numerical methods struggle with high-dimensional problems due to the curse of dimensionality, leading to exponential increases in computational cost and complexity. For strongly nonlinear problems, they often face convergence issues and numerical instability, especially when coupled equations involve multi-scale interactions or solutions with sharp gradients and localized structures like solitons. These limitations have motivated the development of alternative approaches, such as PINNs, to overcome the shortcomings of traditional methods in modern applications. In this paper, an enhanced PINN methodology is introduced for the prediction of soliton structures for (2+ 1)-dimensional CNLSE (1). To the best of our knowledge, the application of PINNs to this class of equations has not been reported before. The proposed approach leverages a neuron-wise locally adaptive activation function to accelerate the convergence rate and improve the accuracy of standard PINN [50]. By combining the proposed approach with the residual-based adaptive refinement strategy (RAR) [51], we demonstrate that each residual loss constructed from the model Eq. (1) requires a different distribution of residual points within the computational domain. The performance and efficiency of the proposed framework will be evaluated in predicting various soliton dynamics through the lens of the $\mathbb{L}^{2}$ relative error. By considering different forms of dispersion and nonlinearity coefficients, vector parabolic dark one-soliton, vector m-shaped anti-dark one-soliton, and dark two-soliton are predicted. We also apply the proposed framework to the parameter discovery of the model Eq. (1). First, we apply the introduced PINN, equipped with a single neural network, for constant parameter discovery under different noise intensities. This PINN model not only predicts the solution of the model equation but also estimates the constant parameter as another trainable network parameter during training. Second, we employ a dual-network strategy for the proposed PINN scheme, where one neural network predicts the solution dynamics, while the other captures the behavior of the dispersion and nonlinearity variable coefficients. The obtained results indicate that the proposed framework is promising for exploring waveform propagation in the optical fiber communications landscape. The main contributions of this study are as follows:
  • 1.
    Introduction of an enhanced PINN methodology for predicting soliton structures governed by the (2+ 1)-dimensional VC-CNLSE (1) for the first time.
  • 2.
    Utilization of a neuron-wise locally adaptive activation function to accelerate convergence rate and improve the accuracy of standard PINN models.
  • 3.
    Integration of the residual-based adaptive refinement (RAR) strategy with the proposed PINN approach to allocate residual points optimally for different residual losses derived from the model equation.
  • 4.
    Estimation of constant parameters using a single neural network under various noise intensities.
  • 5.
    Discovery of variable coefficients using a dual-network strategy.
@@
This is how the remainder of the paper is organized. In Sect. 2, we introduce the general framework of PINN algorithms for the general (2+ 1)-dimensional variable coefficient coupled nonlinear equations and discuss the related mechanisms. We also introduce our enhanced PINN framework in this section. We investigate the application of the proposed enhanced PINN algorithm for data-driven vector dark and anti-dark one- and two-soliton solutions to the (2+ 1)-dimensional VC-CNLSE. We illustratively show that our proposed algorithm accurately captures the solitonic dynamics. We also perform data-driven parameter discovery for the system of model equations to estimate the value of constant coefficients as well as the variable dispersion and nonlinearity coefficients. A conclusion is made in Sec. 4, where we summarize the effectiveness of the enhanced PINN framework and discuss its potential.